OpenStudy (anonymous):

Barbara draws pens randomly from a box containing 5 pens of the same shape and size. There is 1 green pen, 3 red pens, and 1 blue pen. She draws 1 red pen and then another red pen without replacing the first one. Find the probability of drawing 1 red pen followed by another red pen, and show the equation used.

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

(3/5)*100

OpenStudy (anonymous):

60

OpenStudy (anonymous):

2/4*100?

OpenStudy (anonymous):

Which is 50. So (60/100) x (50/100)

OpenStudy (anonymous):

.3

OpenStudy (anonymous):

Yes! And your result is 30%. I don't think it was very clear but did you understand a little bit?

OpenStudy (anonymous):

a little. am just not sure how to put it in essay form

OpenStudy (anonymous):

@jim_thompson5910 am not sure how to explain my answer for this question

OpenStudy (anonymous):

probability = favorable cases / possible cases = 60/100 × 50/100 = (60 × 50)/(100×100) =3000/10000 = 0.3 = 30

OpenStudy (anonymous):

should I leave it as 30 or 30%

OpenStudy (anonymous):

As 30% that's your total. Because it's 50% and 60%

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Your welcome